A cube of a car transmission cube of w is equal to 0.3 m, receiving 150 Horsepower from an engine (1cell =746 w). If the transmission efficiency is 0.93, and the air surrounding temperature is 30°C, and h= 200w/m2K. Calculate surface temperature of the transmission box, if the radiation is not calculated
1
Expert's answer
2020-04-06T10:17:05-0400


From Newton’s law of cooling,


"q=hA_s=(T_s-T_\infty)=6hW^2(T_s-T_\infty)"

The heat rate is


"q=P_i-P_o=P_i(1-\eta)=()()()=7833\ W"

Thus,


"T_s=T_\infty+\frac{q}{6hW^2}=30+\frac{7833}{6(200)(0.3)^2}=102.5\degree C"